Use the following information for the Exercises below [The following information applies to the questions displayed below.] The following financial statements and additional information are reported IKIBAN INC Comparative Balance Sheets June 30, 2018 and 2017 2018 2017 Assets $ 87,500 44,000 51,000 86,500 5,400 186,900 115,000 as Accounts receivable, net Inventory 65,000 63,800 4,400 220,700 124,000 (27,000) Prepaid expenses Total current assets Equipment Accum. depreciation-Equipment Total assets (9,000) $317,700 292,900 Liabilities and Equity Accounts payable $ 25,000 6,000 3,400 34,400 30,000 15,000 3,800 48,800 60,000 108,800 Wages payable Income taxes payable Total current liabilities Notes payable (long term) 30,000 Do Total 1iabilities Equity Common stock, $5 par value Retained earnigs 64,400 160,000 24,100 $317,700 292,900 220,000 33,300 Total liabilities and equity IKIBAN INC. Income Statement For Year Ended June 30, 2018 Sales Cost of goods sold Gross profit Operating expenses $678,000 411,000 267,000 $58,600 67,000 Depreciation expense Other expenses Total operating expenses 125,608 141,400 Other gains (losses) Gain on sale of equipment Income before taxes Income taxes expense Net income 2,000 143,400 43,890 $99,510Additional Information a. A $30,000 note payable is retired at its $30,000 carrying (book) value in exchange for cash. b. The only changes affecting retained earnings are net income and cash dividends paid. c. New equipment is acquired for $57,600 cash. d. Received cash for the sale of equipment that had cost $48,600, yielding a $2,000 gain. e. Prepaid Expenses and Wages Payable relate to Other Expenses on the income statement f. All purchases and sales of inventory are on credit Exercise 12.11 Indirect: Preparing statement of cash flows (part 2) LO PI, P2, P3. Al 2) Compute the companys cash flow on total assets ratio for its fiscal year 2018. Choose Numerator: Choose Denominator: Cash Flow on Total Assets Ratio Operating cash flows / A Average total assets Cash flow on total assets ratio
Answer #1

IKIBAN Inc. Company Statement of Cash Flows Indirect Method For the year ended June 30, 2018 Cash flows from operating activities Net income $99,510 Adjustments to reconcile net income to operating cash flow: Depreciation expense Gain on sale of equipment Increase in Accounts receivable, net (14,000) Decrease in Inventorv Decrease in Prepaid expenses Decrease in Accounts payable Decrease in Wages payable Decrease in Income taxes payable Net cash provided by operating activities $58,600 (2,000) 22,700 1,000 (5,000) (9,000) 400 51.900 151,410 Cash flows from investing activities Cash received sale of equipment Cash paid for equipment Net cash used by investing activities $10,000 (57.600) (47,600) Cash flows from financing activities Cash received from stock issuance Cash paid to retire notes Cash paid for dividends Net cash used by financing activities Net increase in cash Cash balance at June 30, 2017 Cash balance at June 30, 2018 $60,000 (30,000) 90,310 (60.310 $43,500 44,000 $87.500

Cash flow on total assets ratio

Operating cash flows = 151,410

Average total assets = (313,700 + 292,900)/2 = 303,300

Cash flow on total assets ratio = 151,410 / 303,300 = 49.92%
Answer #2
Average total assets = (317,700+292,00) / 2 = 305,300